Find the dimension and a basis for ๐‘Š = {(๐‘ฅ, ๐‘ฆ, ๐‘ง, ๐‘ค,๐‘ก): ๐‘ฅ + ๐‘ฆ + ๐‘ง + ๐‘ค + ๐‘ก = 0, ๐‘ฅ โˆ’ ๐‘ฆ + ๐‘ง โˆ’ ๐‘ค + ๐‘ก = 0}


1
Expert's answer
2022-03-04T09:34:11-0500

"W" is a subspace of "\mathbb{R}^5". From equations "x+y+z+w+t=0,x-y+z-w+t=0" we find: "t=w-z+y-x" . Substitute it in the first equation and get: "x+y+z+w+w-z+y-x=2y+2w=0" . Thus, we get: "t=-z-x", "w=-y." Thus, the space is 3 dimensional. "x,y,z" are independent coordinates and coordinates "t" and "w" can be expressed via "x,y,z" . To find the basis we take the standard basis from "\mathbb{R}^3": "x_1=1,y_1=0,z_1=0", "x_2=0,y_2=1,z_1=0" and "x_3=0,y_3=0,z_3=1" and extend it to "\mathbb{R}^5" by using equations for "t" and "w". We receive 3 vectors: "e_1=(1,0,0,0,-1), e_2=(0,1,0,-1,0)" and "e_3=(0,0,1,0,-1)". This is a basis of "W".